Find the truth value of the statements using the given domain of discourse.

1.Suppose P(x) is the predicate x + 2 =2x and the domain is a set of {1,2,3}.

2.If P(x) is the predicate x2-3x-4=0 and the domain is a set of {1,2,3,4}.



1
Expert's answer
2022-05-02T09:40:11-0400

1. "1+2=3 \not =2"

So it is false

"2+2=4=4"

True

"3+2=5 \not=6"

False

2."1^2-3x1-4=-6 \not =0"

So it is false

"2^2-3x2-4=-6 \not =0"

False

"3^2-3x3-4=-4 \not=0"

False

"4^2-3x4-4=0=0"

True